OpenStudy (anonymous):

Write without exponents. 4 raised to –2 power A) A) –8 B) B) 1/16 C) C) –16 D) D) 1/8

OpenStudy (anonymous):

no with a negative exponent you divide

OpenStudy (anonymous):

\[4^{-2}= \frac{1}{4^2} = \frac{1}{16}\]

OpenStudy (magbak):

It is B

OpenStudy (anonymous):

^